how do you solve this substitution
x+2y=2
5x -3y=-29

Answer by checkley77(12844)

x+2y=2
x=2-2y
Now replace x with (2-2y) in the second equation & solve for y.
5x-3y=-29
5(2-2y-3y=-29
10-10y-3y=-29
-13y=-29-10
-13y=-39
y=-39/-13
y=3 ans.
x=2-2*3
x=2-6
x=-4 ans.
Proof:
5*-4-3*3=-29
-20-9=-29
-29=-29